OpenStudy (anonymous):

If x = y, what does 4(x-y) + 7 =? 4x - 4y + 7 7 4xy + 7 11x - 11y

OpenStudy (anonymous):

@Jadeishere

OpenStudy (anonymous):

@ganeshie8

OpenStudy (jadeishere):

if x = y x- y = 0 4 * 0 = 0 7 = 7 your answer is 7 :)

OpenStudy (anonymous):

Can you please go into more details on how you got that answer please ? but thank you

OpenStudy (jadeishere):

I don't know how to explain it but I can try >.< so it gave you 'x=y' which means if you subtracted y from x, it'd be zero, since they're the same number, and you ALWAYs solve the parentheses first. (PEMDAS) any number multiplied by 0 will be equal to 0 so 4(0) or 4* 0 or 4 x 0 will equal 0 then, all there is left is going to be; 0 + 7 = ? the absolute value of |7|=? is 7 therefore your answer would be 7 >.<